OpenStudy (anonymous):

Given the exponential function f(x)=b^x, show the following: f(u+v)=f(u) * f(v)

jimthompson5910 (jim_thompson5910):

hint: \[\Large a^{b+c} = a^b*a^c\]

OpenStudy (anonymous):

\[F ^{u+v} ?\]

jimthompson5910 (jim_thompson5910):

what is f(u+v) equal to?

jimthompson5910 (jim_thompson5910):

ignore f(u+v)=f(u) * f(v) for a moment

OpenStudy (anonymous):

im kind of confused. What am I supposed to do with the equation you just gave me?

OpenStudy (aum):

\[ \large f(x) = b^x \\ \large f(\text{apples}) = b^{\text{apples}} \\ \large f(u + v) = ? \]

jimthompson5910 (jim_thompson5910):

what is f(x)?

OpenStudy (anonymous):

b^x?

jimthompson5910 (jim_thompson5910):

what happens when you replace every copy of x with u+v

OpenStudy (aum):

If f(x) = b^x then, for example, f(2) = b^2. (We replace wherever there is 'x' with 2) f(u+v) = ?

OpenStudy (anonymous):

b^uv?

OpenStudy (anonymous):

or b^u+b^v

jimthompson5910 (jim_thompson5910):